OpenStudy (yacoub1993):

What is the solution of the system? Use a graph. y = –3x + 2 y = –3x + 4 A. no solutions B. infinitely many solutions C. (0, 4) D. (0, 2)

ganeshie8 (ganeshie8):

graph it and see ? :)

ganeshie8 (ganeshie8):

to graph the lines, you may enter the equations below :- https://www.desmos.com/calculator

OpenStudy (yacoub1993):

is the correct answer C @ganeshie8

OpenStudy (yttrium):

@yacoub, why do you say that C is the answer?

OpenStudy (yacoub1993):

just imagination

OpenStudy (yttrium):

Did you already graph it as ganeshie8 said?

OpenStudy (yacoub1993):

it was graphed with with the choices @Yttrium

OpenStudy (yttrium):

Ahhhh. But only one among those is correct. Anyway, did you found any intersection between those lines?

OpenStudy (yacoub1993):

i thought (0, 4)

OpenStudy (yacoub1993):

not exactly sure

OpenStudy (yttrium):

When you substitute (0,4) 4 = (0) + 2 and 4 = (0) + 4 which is 4 = 2 and 4 = 4 With that, it simply means that it is not the solution of the system.

OpenStudy (yacoub1993):

ohh wahts the solutio

OpenStudy (yttrium):

If you're going to find the real graph of those lines. You will find it no solution at all because they haven't have any intersections. Btw, you know how to graph lines? (I hope so)